Reference Manual

Generated by Doxygen 1.4.1

Fri Jan 19 17:28:44 2007

Contents

1 FINUM (S Crépey, Evry University) 1 2 Namespace Index 3

2.1 Namespace List . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Namespace Documentation 5

3.1 finum Namespace Reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Chapter 1

FINUM (S Crépey, Evry University)

A dll of C/C++ basic financial functions for pedagogical purposes

2  FINUM (S Crépey, Evry University)

Chapter 2

Namespace Index

2.1 Namespace List

Here is a list of all documented namespaces with brief descriptions: finum (Library Functions ) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

4  Namespace Index

Chapter 3

Namespace Documentation

3.1 finum Namespace Reference

Library Functions.

Functions

• double Inverse_erf_Moro (double x)

numerical inverse of the Gaussian c.d.f.

• void tsb2_seq (int ∗n, float x[ ])

Sobol sequence in dimension 1 to 160.

• int Call_BlackScholes_73 (double s, double Maturity, double Strike, double r, double

divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma)

BS call formula.

• int Put_BlackScholes_73 (double s, double Maturity, double Strike, double r, double

divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma)

BS put formula.

• int ImpliedVol (double price, double s0, double Maturity, double Strike, double eqr, double

eqdiv, double sigma_ref, double ∗volimpl, double ∗impliedDelta, double ∗impliedGamma)

BS implied volatility solver (dichotomy).

• int TreePriceNone (double s0, Vanilla Type, double Maturity, double Strike, double r, double divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma, int nTime, double lambda)

BS vanilla tree pricer.

• int ImplPriceNone (double s0, Vanilla Type, double Maturity, double Strike, double r, double divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma, int nTime, int nSpace)

BS vanilla implicit pricer.

• int CNPriceNone (double s0, Vanilla Type, double Maturity, double Strike, double r, double divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma, int nTime, int nSpace)

BS vanilla Crank-Nicholson pricer.

• int TreePriceUpOut (double s0, Vanilla Type, double Maturity, double Strike, double barrier, double rebate, double r, double divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma, int nTime, double lambda)

BS up-and-out barrier tree pricer.

• int ImplPriceUpOut (double s0, Vanilla Type, double Maturity, double Strike, double barrier, double rebate, double r, double divid, double sigma, double ∗ptprice, double ∗ptdelta, double ∗ptgamma, int nTime, int nSpace)

BS up-and-out barrier implicit pricer.

• int main ()

Tests (for console compilation).

3.1.1 Detailed Description

Library Functions.

Index

finum, 5